# Factorized Video Generation: Decoupling Scene Construction and Temporal Synthesis in Text-to-Video Diffusion Models

### v1

@@ 6 Ablation Studies
To better understand the components contributing to the effectiveness of factorized video generation, we conduct two ablations. First, we analyze the impact of the LLM-based prompt rewriting for anchor image generation. Second, we evaluate the model’s sensitivity to different anchor images.
@@ 6.2 Sensitivity to the Choice of Anchor Image
While our main experiments use a single preferred anchor image, mimicking a human-in-the-loop workflow, we also evaluate how sensitive the model is to the specific anchor provided. For Wan 5B, we generate five different anchor images per prompt using our LLM-refined first-frame description and assess the resulting videos across all VBench 2.0 categories. As shown in Tab. 5, the variation across different image anchors is extremely small: all categories have standard deviations below 0.02. This indicates that the slight variations resulting from the stochastic generation of anchor images affect mostly appearance and not the model’s compositional performance.
To contextualize this robustness, Tab. 5 compares the variability of our factorized model across anchors with the variability of the original T2V model across random seeds. On average, T2V is 9× more variable with respect to seed than with respect to image anchors. This indicates the effectiveness of our factorized approach.
Table 5: Standard deviation across anchor images (Factorized Wan5B) vs. across seeds (Wan5B). Lower values indicate greater stability.
@@ 7 Discussion
Our results suggest that improved grounding, rather than increased capacity alone, may be equally important. Recent advances in T2V diffusion have relied heavily on increasing model size and training data, yet even large models often struggle to infer a coherent initial scene from text alone. This contrasts with image diffusion, where scaling is relatively straightforward; in video models, each architectural improvement must operate over an additional temporal dimension, making scaling substantially more resource intensive. Our findings indicate that improved grounding can complement scale by addressing a different bottleneck: establishing the correct scene before motion synthesis begins. By factorizing video generation into scene composition and temporal modeling, we mitigate several common failure modes without requiring substantially larger models. We view this as a complementary design principle that can guide future architectures toward more reliable and structured video synthesis.
Implications for T2V evaluation.
@@
The effectiveness of anchor-based generation also highlights a potential gap in current evaluation practices. Benchmarks such as VBench2.0 and T2V-CompBench reward accurate scene composition and prompt adherence, yet we find that standard T2V models often underperform relative to a simple factorized baseline. This suggests that future evaluations may benefit from including anchor-conditioned baselines or I2V+text baselines to better disentangle scene construction from true temporal modeling. Ideally, to reflect improvements in motion understanding and generation of T2V models rather than their improvements in static composition, it would be valuable to evaluate their factorized version.
Performance on Human Fidelity.
@@
As shown in Tab. 3, both prompt upsampling and visual anchoring tend to reduce performance on the Human Fidelity category. This aggregate score combines three sub-metrics: human anatomy, human clothes, and human identity. As detailed in the full results in Tab. 6 (Appendix), the decline is driven almost entirely by the human identity component, while the other two components remain stable. Through manual inspection, we observe that all model variants, regardless of anchoring, struggle to preserve identity especially across different actions. While our primary focus is improving compositional and reasoning capabilities, identity preservation is a distinct and challenging problem in its own. Addressing this limitation, potentially through explicit identity conditioning or persistent character representations, remains an important direction for future work.
On the diversity–correctness trade-off.
@@
Diversity is essential for applications such as synthetic data generation and counterfactuals, where one must produce multiple distinct videos without violating the intended scenario. Yet current diversity metrics rely on VGG feature differences, which primarily capture scene appearance. As shown in our performance ablation in Sec. 6.2, variability across seeds can meaningfully affect correctness, meaning that these metrics may interpret clear compositional failures or inconsistent layouts as “diversity”. Our factorized pipeline reduces prompt-adherence failures, as evidenced by its stronger benchmark scores, but while anchoring improves correctness, it does not guarantee diversity under current metrics. Conversely, standard T2V achieves higher numerical diversity but explores a broader but less controlled generative space, sometimes producing unfaithful or incorrect outcomes. We argue that the introduction of faithful diversity metric ensuring not only that the outputs vary but also that they respect certain constraints would be highly valuable. We also suggest that true video diversity should reflect two distinct diversity dimensions. First, the appearance diversity of the frames, which is captured with current metrics. Second, the motion diversity under given conditions. We argue that a possible principled way to evaluate motion diversity is to quantify the variability in a factorized model’s outputs when the first frame is kept constant, isolating motion from scene composition.
@@ 8 Conclusion
We proposed Factorized Video Generation, a simple pipeline that grounds the first frame before synthesizing motion, addressing a key failure mode of current T2V models. This factorization yields strong gains on compositional and reasoning benchmarks, with small models outperforming larger baselines and maintaining accuracy at one-third the sampling steps. While challenges such as identity preservation remain, our results show that explicitly separating scene construction from temporal modeling is a powerful and practical direction for future video generation research.
@@ 9 Acknowledgments
Appendix A FLops Analysis
In this section, we give a more detailed study of the computation required by each model. The factorized version required a diffusion model for image generation and one for video generation. We report the Flops for a single call for each model. During sampling, the models are called multiple times for inference. As a video model, we report the Flops for Wan 2.2 with 5 billion parameters. For the image model, we report the Flops for Qwen-Image with 20 billion parameters. A single call of Wan 2.2 represents $117\times 10^{3}$ GFLOPS. To generate a video, we use 50 inference steps and classifier-free guidance. As a result, the total cost of generating a video is $117\times 10^{5}$ GFLOPS. A single call of Qwen-Image represents $10.7$ GFlOPS. To generate an image, we use 50 inference steps and classifier-free guidance. As a result, generating an image costs about $10.7\times 10^{2}$ GFLOPS. Finally, we use the factorized approach to reduce the number of inference steps from 50 to 15 for video generation. The total cost is $35.1\times 10^{5}$ GFLOPS.
Appendix B Detailed Experiments Results
B.1 VBench 2.0
Table 6 show the detailed metric results for VBench2.0 averaged across five seeds for Wan2.2 5B, Wan2.2 14B and CogVideo1.5 5B. Across the full set of 17 VBench2.0 metrics, we observe consistent trends that mirror the aggregated results in Tab. 3. First, factorized T2V improves nearly every metric relative to both T2V and upsampled T2V for all three model sizes (5B, 14B, and 5B CogVideoX). Improvements are especially significant in categories reflecting reasoning-driven scene understanding such as Composition, Instance Preservation, Motion Order Understanding, Dynamic Attribute and Dynamic Spatial Relationship further supporting our claim that anchoring mitigates the dominant failure mode of T2V: incorrect scene construction.
@@ Appendix B Detailed Experiments Results
Third, consistent with our observations in Tab. 3, anchoring does not improve performance on Human Identity. The identity-related metric drops even when using upsampled prompts, indicating that this limitation stems from the underlying generative model.
@@
Factorized T2V 0.7420 0.8070 0.7587 0.7970 0.1859 0.4655 0.0150 0.6761
B.2 Robustness to Reduced Sampling Steps
Table 7 and the full metric values reported in Table 8 show a clear difference in robustness of each model to a reduction in number of sampling steps. The text-only T2V model’s performance drops at 30 steps (–3.5% on average) and collapses at 15 steps (–17%), with large drop in categories such as Action Binding and Object Interaction. The upsampled T2V variant follows a similar pattern. Due to the fact that it begins with a higher 50-step baseline, its average performance decreases by –10.2% at 30 steps and –22.0% at 15 steps, with some metrics falling by over –35%. On the other hand, the factorized model remains remarkably stable. Even when the sampling steps are reduced to 30 or 15, its average score changes by only +1.22% and +0.26%, respectively. These results indicate that visual anchoring significantly improves robustness to aggressive step reduction. Thus, once the initial scene is grounded, generation relies far less on long denoising steps.
@@ B.2 Robustness to Reduced Sampling Steps
Factorized T2V 15 0.26 -2.27 0.18 -1.66 0.03 2.96 -0.50 0.26
Factorized T2V 30 0.02 0.68 0.16 -0.35 1.65 3.66 1.23 1.22
Table 8: Average WAN-5B results across five seeds for T2V, upsampled T2V, and our factorized T2V at different sampling steps. Our method remains stable even at 15 steps, whereas both T2V variants degrade sharply as steps decrease.
@@
Factorized T2V 15 0.9399 0.3836 0.6606 0.2799 0.8214 0.8953 0.7512 0.6760
Factorized T2V 30 0.9377 0.3951 0.6604 0.2836 0.8347 0.9013 0.7642 0.6824
Factorized T2V 50 0.9375 0.3925 0.6594 0.2846 0.8212 0.8695 0.7550 0.6742
Appendix C Qualitative Results
@@
Figure 5, Fig. 6, and Fig. 7 provides additional examples illustrating the visual differences between standard T2V generation and our factorized approach for Wan14B, Wan5B and CogVideo5B respectively across a wide variety of scenes. Across diverse prompts, the anchored model produces scenes with more accurate semantic alignment with the described situation. On the other hand, text-only T2V baselines often hallucinate or misinterpret spatial relations, or produce semantically incorrect first frames. These examples complement our quantitative results and highlight how grounding the visual anchor substantially improves scene construction.
Figure 5: Overall qualitative results comparing our factorized method against T2V methods. Visual examples from Wan14B showing that factorized T2V produces more coherent scene layouts and semantically aligned compositions than text-only T2V methods..
Figure 6: Overall qualitative results comparing our factorized method against T2V methods. Visual examples from Wan5B showing that factorized T2V produces more coherent scene layouts and semantically aligned compositions than text-only T2V methods..
Figure 7: Overall qualitative results comparing our factorized method against T2V methods. Visual examples from CogVideo showing that factorized T2V produces more coherent scene layouts and semantically aligned compositions than text-only T2V methods.
@@
Figure 8 compares anchor images generated directly from the raw video prompt with those produced using the LLM-refined first-frame description. The refined prompts generate anchors that more accurately capture the intended initial configuration of objects, spatial layout, and contextual cues. Videos produced from these anchors exhibit stronger compositional fidelity and temporal coherence. This qualitative trend aligns with the quantitative improvements reported in Sec. 6.1.
@@
Figure 9 illustrates how decreasing the number of diffusion steps affects visual quality. For text-only T2V models, fewer steps lead to degraded scene structure throughout the entire video. The factorized model preserves scene integrity even at 15 sampling steps. These results visually reinforce the stability trends observed quantitatively in Fig. 5, demonstrating that anchoring makes denoising far more robust to aggressive sampling reduction.
Figure 9: Qualitative sensitivity to sampling steps. Factorized T2V maintains coherent scene structure even at 15 steps, whereas text-only T2V collapses as sampling decreases.
@@
Figure 10 illustrates typical failure modes of prompt upsampling. Although upsampling produces more detailed text descriptions, it does not correct the underlying compositional weaknesses of T2V models as they seem to lead to similar problems in terms of semantic composition. Our factorized model resolves these cases by grounding generation in a concrete visual anchor. These examples highlight that text refinement alone cannot substitute for explicit scene grounding.
Figure 10: Prompt upsampling vs. factorized T2V. Upsampled T2V improves textual detail but does not fix core compositional failures. Our factorized model successfully reconstructs the correct scene structure across the same prompts.

### v2

@@ 6 Ablation Studies
To better understand the components contributing to the effectiveness of our framework, we conduct three ablations. First, we study the flexibility of anchor placement by evaluating anchoring at different temporal positions. Second, we analyze the impact of LLM-based prompt rewriting for anchor image generation. Third, we evaluate the model’s sensitivity to different anchor images.
6.1 Flexible Anchor Position
Anchor T2VComp. VB2
First 0.661 0.646
Middle 0.670 0.631
Table 5: AVG Wan14B with different anchor positions.
We study whether AVG depends on first-frame anchoring, or whether anchors can be placed at other temporal positions. To test this, we modify the reasoning stage to predict the middle-frame scene state and use this temporally aligned description for composition and anchoring. This is enabled by our anchor-grounding finetuning, which allows a clean injected frame to serve as a valid anchor at arbitrary positions. As shown in Table 5, middle-frame anchoring remains comparable to first-frame anchoring. This suggests that the exact anchor location is not the key factor; what matters is producing the correct visual state at the chosen temporal position. Once the scene is properly grounded, temporal synthesis becomes easier, as the model can focus on motion rather than jointly solving scene construction and animation. Overall, this highlights AVG as a flexible pipeline that can incorporate anchors at different temporal locations.
@@ 6.3 Sensitivity to the Choice of Anchor Image
While our main experiments use a single preferred anchor image, mimicking a human-in-the-loop workflow, we also evaluate how sensitive the model is to the specific anchor provided. For Wan 5B, we generate five different anchor images per prompt using our LLM-refined first-frame description and assess the resulting videos across all VBench 2.0 categories. As shown in Tab.˜6, the variation across different image anchors is extremely small: all categories have standard deviations below 0.02. This indicates that the slight variations resulting from the stochastic generation of anchor images affect mostly appearance and not the model’s compositional performance. To contextualize this robustness, Tab.˜6 compares the variability of our anchored model across anchors with the variability of the original T2V model across random seeds. On average, T2V is 9× more variable with respect to seed than with respect to image anchors. This indicates the effectiveness of our anchored approach.
Table 6: Standard deviation across anchor images (Anchored Wan5B) vs. across seeds (Wan5B). Lower values indicate greater stability.
@@ 7 Discussion
Anchoring vs. model scaling. Our results suggest that improved grounding, rather than increased capacity alone, may be equally important for T2V generation. Recent advances in T2V diffusion have relied heavily on scaling model size and training data, yet even large models often struggle to infer a coherent initial scene from text alone. This contrasts with image diffusion, where scaling is relatively straightforward; in video models, each architectural improvement must operate over an additional temporal dimension, making scaling substantially more resource intensive. Our findings indicate that stronger visual grounding can complement scale by addressing a different bottleneck: establishing the correct scene before motion synthesis. By separating scene composition from temporal modeling through a visual anchor, we mitigate several common failure modes without requiring substantially larger models. We view this as a complementary design principle for building more reliable and structured video models.
@@
Implications for T2V evaluation. The effectiveness of visual anchoring also highlights a potential gap in current evaluation practice. Benchmarks such as VBench2.0 and T2V-CompBench reward accurate scene composition and prompt adherence, yet we find that standard T2V models often underperform relative to a simple anchored baseline. This suggests that future evaluations may benefit from including anchor-conditioned or native I2V+text baselines to better disentangle scene construction from true temporal modeling. More broadly, if the goal is to measure progress in motion understanding rather than improvements in static composition alone, it may also be valuable to evaluate anchored counterparts of T2V models under the same backbone.
@@
Performance on Human Fidelity. As shown in Tab.˜3, both prompt upsampling and visual anchoring tend to reduce performance on the Human Fidelity category. This aggregate score combines three sub-metrics: human anatomy, human clothes, and human identity. As detailed in the full results in Tab.˜7 (Appendix), the decline is driven almost entirely by the human identity component, while the other two components remain stable. Through manual inspection, we observe that all model variants, regardless of anchoring, struggle to preserve identity especially across different actions. While our primary focus is improving compositional and reasoning capabilities, identity preservation is a distinct and challenging problem in its own. Addressing this limitation, potentially through explicit identity conditioning or persistent character representations, remains an important direction for future work.
@@
On the diversity–correctness trade-off. Diversity is important for applications such as synthetic data generation and counterfactuals, where multiple distinct videos must be produced without violating the intended scenario. However, current diversity metrics are based on VGG feature differences and mainly capture scene appearance. As shown in our performance ablation in Sec.˜6.3, variation across seeds can significantly affect correctness, so these metrics may count compositional failures or inconsistent layouts as “diversity.” Our anchored pipeline reduces prompt-adherence failures, as reflected in its stronger benchmark scores, but although anchoring improves correctness, it does not guarantee diversity under existing metrics. Conversely, standard T2V attains higher numerical diversity by exploring a broader but less controlled generative space, which can also yield unfaithful or incorrect outputs. We therefore argue that a faithful diversity metric should measure not only whether outputs vary, but also whether they satisfy the intended constraints. More specifically, true video diversity should capture two dimensions: (1) appearance diversity across frames, which current metrics partially measure, and (2) motion diversity under fixed conditions. A principled way to evaluate motion diversity may be to measure variability in an anchored model’s outputs while keeping the first frame fixed, thereby isolating motion from scene composition.
@@ 8 Conclusion
We proposed a simple anchored generation framework, a simple pipeline that grounds the first frame before synthesizing motion, addressing a key failure mode of current T2V models. This explicit visual anchoring yields strong gains on compositional and reasoning benchmarks, with small models outperforming larger baselines and maintaining accuracy at one-third the sampling steps. While challenges such as identity preservation remain, our results show that explicitly separating scene construction from temporal modeling is a powerful and practical direction for future video generation research.
@@ 9 Acknowledgments
Appendix 0.A FLops Analysis
In this section, we give a more detailed study of the computation required by each model. The anchoredd version required a diffusion model for image generation and one for video generation. We report the Flops for a single call for each model. During sampling, the models are called multiple times for inference. As a video model, we report the Flops for Wan 2.2 with 5 billion parameters. For the image model, we report the Flops for Qwen-Image with 20 billion parameters. A single call of Wan 2.2 represents $117\times 10^{3}$ GFLOPS. To generate a video, we use 50 inference steps and classifier-free guidance. As a result, the total cost of generating a video is $117\times 10^{5}$ GFLOPS. A single call of Qwen-Image represents $10.7$ GFlOPS. To generate an image, we use 50 inference steps and classifier-free guidance. As a result, generating an image costs about $10.7\times 10^{2}$ GFLOPS. Finally, we use the anchored approach to reduce the number of inference steps from 50 to 15 for video generation. The total cost is $35.1\times 10^{5}$ GFLOPS.
Appendix 0.B Detailed Experiments Results
0.B.1 VBench 2.0
Table˜7 show the detailed metric results for VBench2.0 averaged across five seeds for Wan2.2 5B, Wan2.2 14B and CogVideo1.5 5B. Across the full set of 17 VBench2.0 metrics, we observe consistent trends that mirror the aggregated results in Tab.˜3. First, anchored T2V improves nearly every metric relative to both T2V and upsampled T2V for all three model sizes (5B, 14B, and 5B CogVideoX). Improvements are especially significant in categories reflecting reasoning-driven scene understanding such as Composition, Instance Preservation, Motion Order Understanding, Dynamic Attribute and Dynamic Spatial Relationship further supporting our claim that anchoring mitigates the dominant failure mode of T2V: incorrect scene construction.
@@ Appendix 0.B Detailed Experiments Results
Third, consistent with our observations in Tab.˜3, anchoring does not improve performance on Human Identity. The identity-related metric drops even when using upsampled prompts, indicating that this limitation stems from the underlying generative model.
@@
Anchored T2V 0.7420 0.8070 0.7587 0.7970 0.1859 0.4655 0.0150 0.6761
0.B.2 Robustness to Reduced Sampling Steps
Table 8 and the full metric values reported in Table 9 show a clear difference in robustness of each model to a reduction in number of sampling steps. The text-only T2V model’s performance drops at 30 steps (–3.5% on average) and collapses at 15 steps (–17%), with large drop in categories such as Action Binding and Object Interaction. The upsampled T2V variant follows a similar pattern. Due to the fact that it begins with a higher 50-step baseline, its average performance decreases by –10.2% at 30 steps and –22.0% at 15 steps, with some metrics falling by over –35%. On the other hand, the anchored model remains remarkably stable. Even when the sampling steps are reduced to 30 or 15, its average score changes by only +1.22% and +0.26%, respectively. These results indicate that visual anchoring significantly improves robustness to aggressive step reduction. Thus, once the initial scene is grounded, generation relies far less on long denoising steps.
@@ 0.B.2 Robustness to Reduced Sampling Steps
Anchored T2V 15 0.26 -2.27 0.18 -1.66 0.03 2.96 -0.50 0.26
Anchored T2V 30 0.02 0.68 0.16 -0.35 1.65 3.66 1.23 1.22
Table 9: Average WAN-5B results across five seeds for T2V, upsampled T2V, and our anchored T2V at different sampling steps. Our method remains stable even at 15 steps, whereas both T2V variants degrade sharply as steps decrease.
@@
Anchored T2V 15 0.9399 0.3836 0.6606 0.2799 0.8214 0.8953 0.7512 0.6760
Anchored T2V 30 0.9377 0.3951 0.6604 0.2836 0.8347 0.9013 0.7642 0.6824
Anchored T2V 50 0.9375 0.3925 0.6594 0.2846 0.8212 0.8695 0.7550 0.6742
Appendix 0.C Qualitative Results
@@
Figure˜5, Fig.˜6, and Fig.˜7 provides additional examples illustrating the visual differences between standard T2V generation and our anchored approach for Wan14B, Wan5B and CogVideo5B respectively across a wide variety of scenes. Across diverse prompts, the anchored model produces scenes with more accurate semantic alignment with the described situation. On the other hand, text-only T2V baselines often hallucinate or misinterpret spatial relations, or produce semantically incorrect first frames. These examples complement our quantitative results and highlight how grounding the visual anchor substantially improves scene construction.
Figure 5: Overall qualitative results comparing our anchored method against T2V methods. Visual examples from Wan14B showing that anchored T2V produces more coherent scene layouts and semantically aligned compositions than text-only T2V methods..
Figure 6: Overall qualitative results comparing our anchored method against T2V methods. Visual examples from Wan5B showing that anchored T2V produces more coherent scene layouts and semantically aligned compositions than text-only T2V methods..
Figure 7: Overall qualitative results comparing our anchored method against T2V methods. Visual examples from CogVideo showing that anchored T2V produces more coherent scene layouts and semantically aligned compositions than text-only T2V methods.
@@
Figure 8 compares anchor images generated directly from the raw video prompt with those produced using the LLM-refined first-frame description. The refined prompts generate anchors that more accurately capture the intended initial configuration of objects, spatial layout, and contextual cues. Videos produced from these anchors exhibit stronger compositional fidelity and temporal coherence. This qualitative trend aligns with the quantitative improvements reported in Sec.˜6.2.
@@
Figure˜9 illustrates how decreasing the number of diffusion steps affects visual quality. For text-only T2V models, fewer steps lead to degraded scene structure throughout the entire video. The anchored model preserves scene integrity even at 15 sampling steps. These results visually reinforce the stability trends observed quantitatively in Fig.˜5, demonstrating that anchoring makes denoising far more robust to aggressive sampling reduction.
Figure 9: Qualitative sensitivity to sampling steps. anchored T2V maintains coherent scene structure even at 15 steps, whereas text-only T2V collapses as sampling decreases.
@@
Figure 10 illustrates typical failure modes of prompt upsampling. Although upsampling produces more detailed text descriptions, it does not correct the underlying compositional weaknesses of T2V models as they seem to lead to similar problems in terms of semantic composition. Our anchored model resolves these cases by grounding generation in a concrete visual anchor. These examples highlight that text refinement alone cannot substitute for explicit scene grounding.
Figure 10: Prompt upsampling vs. anchored T2V. Upsampled T2V improves textual detail but does not fix core compositional failures. Our anchored model successfully reconstructs the correct scene structure across the same prompts.